Following is a list of national income figures for a certain year. All figures are in billions.

Interest and investment income

$
7
Government current purchases of goods and services


32
Taxes less subsidies on factors of production


18
Indirect business taxes less subsidies


6
Personal consumption expenditures


133
Net investment (net capital formation)


37
Capital consumption allowances (depreciation)


23
Net Income of farms and unincorporated businesses


17
Net exports


-8
Profits of corporations and government enterprises before taxes


38
Wages, salaries, and supplementary labour income


108

Calculate GDP using the expenditure method.

GDP=$?? (IN BILLIONS)
1
Expert's answer
2012-01-26T13:06:08-0500
GDP (Y) is a sum of Consumption (C), Investment (I), Government Spending (G) and
Net Exports (X – M). So Y = C + I + G + (X − M). C is Personal consumption
expenditures = 133; I is the sum of Net investment and Depreciation = 37+23 =
60; G is Government current purchases of goods and services = 32; X-M = -8. Thus
GDP = $217 billion.